OpenStudy (anonymous):

A data set consists of the following data points: (3,5), (5,8), (6,13) The slope of the best fit line is 2.5. Find the y-intercept of this line.

OpenStudy (amistre64):

the anchor point is going to be the average of the other 3 add em up and divide by 3

OpenStudy (amistre64):

(3,5), (5,8), (6,13) ------ (14,26)/3 then use the point slope form of a line y - 26/3 = m(x-14/3)

OpenStudy (amistre64):

a little swapping about and we get: y = mx+b; where b is the yint nad equate to: -m*14/3 + 26/3

OpenStudy (anonymous):

Ohk. i think i understand!

OpenStudy (amistre64):

its been awhile, but i do believe thats how you find the "common" point to anchor the slope to its the average of the others

OpenStudy (anonymous):

I got -3

OpenStudy (amistre64):

-2.5*14/3 + 26/3 = -3; me too :)